OpenStudy (anonymous):

If a thimble holds 5 ml of water, how many liters of water does it hold?

OpenStudy (anonymous):

5/1000 = 1/200

OpenStudy (jacob902):

5/1000 = 1/200 Ok, for this you need to know that there are 1000ml in 1L. So 5/1000 = 0.005. The thimble holds 0.005 liters of water.